# Charge Regulation Effect on Nanoparticles Interaction Mediated by Polyelectrolyte

### v1

@@ III Computation of total force between two nanoparticles
We now focus on the analytical form of the total force between two NPs due to bathing PE solution. In the limit of CC, the interaction forces between charged colloids mediated by PEs have already been discussed in detail in Ref. Podgornik et al. (1995). However, we present a pedagogical derivation of these results to ensure that the paper is self-contained. To evaluate the total force acting between two NPs, we introduce a fictitious plane, $S$, of dimension $0.25\ell_{B}\times 80\ell_{B}\times 80\ell_{B}$, positioned at the midpoint between the two NPs (see Fig. 2)  Podgornik et al. (1995); Granfeldt et al. (1991). The system is divided into two sub-volumes, $V_{1}$ and $V_{2}$, corresponding to the left and right halves of the simulation domain, with $S$ acting as the dividing midplane. Given Ref. Podgornik et al. (1995), the total force ($F_{X}$) acting along the radius vector joining the two NPs can be cast as,
$F_{X}=k_{B}T\int_{(S)}d^{2}\mathbf{r}\rho_{1}(\mathbf{r})+\int_{(V_{1})}\int_{(V_{2})}d^{3}\mathbf{r_{1}}d^{3}\mathbf{r_{2}}\rho_{2}(\mathbf{r_{1}},\mathbf{r_{2}})\times\mathbf{f_{x}}(\mathbf{r_{1}},\mathbf{r_{2}})$ (4)
@@
The first term of Eq. 4 represents the momentum exchange between $V_{1}$ and $V_{2}$ due to the movement of ions/particles across the midplane. It has the form of an ideal osmotic pressure Lovett and Baus (1992). The second term involves the microscopic force $\mathbf{f_{x}}(\mathbf{r_{1},r_{2}})$, which acts between a particle at position $\mathbf{r_{1}}\in V_{1}$ and another at $\mathbf{r_{2}}\in V_{2}$ projected along the direction normal to the dividing plane $S$. Here, $\rho_{1}(r)$ refers to the single particle density at the midplane and $\rho_{2}(\mathbf{r_{1}},\mathbf{r_{2}})$ is the two-particle correlation function for the system. This expression for force (Eq. 4) is related to the contact theorem in plan-parallel geometry Lovett and Baus (1992). The microscopic force $\mathbf{f_{x}}(\mathbf{r_{1},r_{2}})$ comprises several contributions: the electrostatic force between the polymer beads, the electrostatic force between the polymer beads and the two NPs, the configurational (elastic) forces arising from bonded polymer segments whose connecting vector intersects the dividing plane $S$.
Figure 2: Schematic of the system for the force calculation. A fictitious plane, $S$, of width $0.25\ell_{B}$ is placed at the midpoint, which divides the system into two halves $V_{1}$ and $V_{2}$. The color scheme is the same as in Fig. 1.
As in Ref. Podgornik et al. (1995) we obtain the total force,
$F_{X}=F_{X}(\text{osm})+F_{X}(\text{cor})+F_{X}(\text{nn})+F_{X}(\text{bridge})$ (5)
@@
where $F_{X}(\text{osm})$ denotes the contribution from the osmotic force due to the momentum exchange mechanism across planes. The force component $F_{X}(\text{cor})$ is due to the correlation effect, and the contribution due to electrostatic interaction between CR NPs is captured by $F_{X}(\text{nn})$. Since the polymer can act as a bridge between NPs when crossing mid-plane $S$, there is an additional term $F_{X}(\text{bridge})$ in the above Eq. 5.
@@
The osmotic force component, $F_{X}(\text{osm})$, is calculated using the relation $F_{X}(\text{osm})=kT<n>$, where $<n>$ is the average number of particles inside the plane $S$. The correlation force component, $F_{X}(\text{cor})$, is obtained by evaluating the electrostatic interactions between (i) all mobile particles in subvolumes $V_{1}$ and $V_{2}$; (ii) all mobile particles in $V_{1}$ and the NP in $V_{2}$; and (iii) the NP in $V_{1}$ and all the mobile particle in $V_{2}$. $F_{X}(\text{nn})$ is direct electrostatic force between two NPs. We have utilized the Particle–Particle Particle–Mesh (PPPM) concept to compute electrostatic forces with a precision of $10^{-3}$. The bridging force, $F_{X}(\text{bridge})$, given by
$F_{X}(\text{bridge})=-K_{F}\left<|X_{i}-X_{i+1}|\right>\times n_{br}$ (6)
@@ IV Results and Discussion
In this section, we present numerical results based on the formulation described above. To elucidate the effects of charge regulation (CR) on the PE-NP interaction and NP-NP interaction, we compare the results of CR and constant charge (CC) simulations performed under the same conditions, as outlined in Sec. II.
@@
We begin with the case of a deionized solution, i.e., with $pIp=pIm=10^{-7}M$. The corresponding system parameters are $pH=7$, $pK_{a}=pK_{b}=5$, and the number of monomers is fixed at $N=40$. Figure 3a shows simulation snapshots of CR and CC systems at a fixed intersurface separation of $D=7\ell_{B}$. Notably, in the CR case, the PE is adsorbed onto the surface of one of the NPs (a localized state). In contrast, under CC conditions at the same separation, PE preferentially forms a bridge between two like-charged NPs, referred to here as the captured state. Within the CC approximation, bridging interactions occur due to the electrostatic attraction between oppositely charged PE chain and the fixed surface charges of NPs, creating a strong and persistent electrostatic attraction for the oppositely charged PE segments, leading to bridging. As a result, the PE becomes partially adsorbed onto both nanoparticles, and the intervening segment forms a bridge that generates an entropic-elastic force between them. These observations are consistent with previous studies Podgornik (2003); Podgornik and Ličer (2006). In contrast, under CR conditions, the surface charges of the nanoparticles dynamically adjust in response to the local electrostatic environment. This flexibility allows one of the nanoparticles to acquire a higher effective surface charge that strongly attracts the entire PE chain, resulting in its complete adsorption. This adaptive charge response suppresses polymer bridging and favors a localized state, thereby minimizing the system’s free energy through a combination of enhanced electrostatic attraction and configurational entropy gain. Further, we find that, under CR conditions, the PE is equally likely to adsorb onto either of the two NPs, reflecting the system’s symmetry.
@@
At equilibrium, the total charge on the PE and the adsorbing NPs is nearly equal in magnitude and opposite in sign for the CR parameters mentioned above, as illustrated in Fig. 3b. Figure 3c presents the total force ($F_{X}$) between the NPs as a function of their intersurface separation. The results show that the PE mediates an attractive interaction between like-charged NPs over a broad range of separations, consistent with previous findings for the CC case Podgornik et al. (1995). In contrast, for the CR case, attractive interactions are observed only at very short intersurface separations and diminish rapidly with increasing distance, becoming weakly repulsive around $D=7\ell_{B}$. This behavior results from the complete adsorption of the PE onto one of the NPs, effectively neutralizing its surface charge. As the separation increases, the dominant contribution to the total force arises from the osmotic pressure of mobile ions and monomers near the midplane, leading to a net repulsive interaction. Notably, charge regulation significantly alters the interaction landscape. As seen in Fig. 3c at $D=15\ell_{B}$, the magnitude of $F_{X}$ in the CR case differs by nearly a factor of six times compared to that in CC. However, at very large separations, the total forces in both CR and CC converge. This is because, in the CC scenario, the PE also eventually undergoes complete adsorption onto an NP.
@@
Figure 3: Charge regulation (CR) effects on NPs–PE interactions at low salt concentrations ($pIp=pIm=7$) with PE bearing acidic groups ($pK_{a}=5$) and NPs coated with basic groups ($pK_{b}=5$). (a) Top: Snapshot of CR simulation at intersurface separation between colloids, $D=7\ell_{B}$. Red and white beads denote charged and neutral acid monomers of the PE, while blue and white sites represent charged and neutral base sites. Bottom: Simulation snapshot of a constant charge (CC) simulation at the same $D$. (b) Average charge on NPs and PE as a function of $D$. (c) Total force ($F_{X}$), between NPs for CR and CC as a function of $D$. (d) Adsorbed monomer fraction ($f_{ad}$), within $2\sigma$ of the colloid surface as a function of $D$. (e) The radius of gyration ($R_{g}$) of the PE as a function of $D$. Data averaged over 7 (CR) and 30 (CC) equilibrated configurations.
@@
A quantitative characterization of the CR effect is provided by examining the fraction of adsorbed monomers, $f_{ad}$, defined as those located within $2\sigma$ of the NP surface (see Fig. 3d). At very small intersurface separations, monomers remain near the NP surfaces due to limited available space, resulting in similar values of $f_{ad}$ and $F_{X}$ for both CR and CC simulations. However, as $D$ increases beyond $7\ell_{B}$, the PE in CR simulations remains fully adsorbed onto one of the NP surfaces. As a consequence, attractive interactions are suppressed, since they originate from monomer stretching across the midplane. In contrast, for the CC case, $f_{ad}$ decreases with increasing $D$, indicating that the PE bridges the two NPs. As separation increases, monomers span a larger distance across the midplane, leading to stronger entropic-elastic attractions. Beyond a critical separation (around $15\ell_{B}$), the PE in CC simulations begins to adsorb onto a single NP surface, reducing the bridging and hence diminishing the attractive force. Even at very large separations, the CR model shows nearly twice the monomer adsorption compared to CC. This suggests that CR enhances the PE’s affinity for the surface, leading to stronger adhesion. This difference is also reflected in the radius of gyration, $R_{g}$, shown in Fig. 3e, which remains consistently higher in CC simulations. The increased $R_{g}$ indicates a more extended polymer conformation due to weaker adsorption and stronger monomer-monomer repulsions. These findings indicate that CR significantly enhances PE adsorption onto NPs, in agreement with recent studies on PE adsorption onto planar surfaces Yuan and Tanaka (2024).
@@
Next, we examine the influence of PE chain length ($N$) on adsorption behavior and the interaggregate force ($F_{X}$). Simulations were performed for $N=30$, $40$, and $50$ while keeping all other parameters identical to those used in Fig. 3. As shown in Fig. 4a, the PE chain length exhibits minimal impact on $F_{X}$ in the CR simulations. This insensitivity can be attributed to the strong adsorption of the PE onto the surface of a single NP, as illustrated in Fig. 4b, which effectively neutralizes the NPs charge and suppresses long-range electrostatic interactions. Interestingly, Fig.4b shows that PE adsorption onto the NP surface monotonically increases with polymer chain length. This behavior is in stark contrast to previous observations for planar surfaces Yuan and Tanaka (2024), where adsorption tends to decrease with increasing chain length due to the entropic penalty associated with flattening long chains onto a surface. In our case, the curved geometry of the NP reduces the conformational entropy loss during adsorption, allowing longer chains to maintain configurational flexibility while still forming multiple contacts with the surface. Additionally, the enhanced total charge of longer PEs increases their electrostatic attraction to the charged NP, further promoting adsorption. The CC model reveals a pronounced dependence of $F_{X}$ on the PE chain length. As depicted in Fig. 4c, longer chains lead to stronger attractive interactions at larger intersurface separations. This enhancement is driven by more effective bridging between the NPs, consistent with the observed decrease in PE adsorption onto individual NP surfaces (see Fig. 4d). Thus, while the interaggregate force in the CR case remains largely unaffected by PE length due to complete adsorption, the CC case exhibits a strong length dependence due to persistent bridging interactions.
@@
Figure 4: Effect of PE chain length on interaggregate force and PE adsorption across different simulation models, under the same conditions as in Fig. 3. (a) Total force ($F_{X}$) between NPs in CR simulation. (b) Adsorbed monomer fraction onto NPs in CR simulations. (c) Total force ($F_{X}$) between NPs in CC simulations. (d) Adsorbed monomer fraction onto NPs in CC simulation.
@@
We now investigate the effect of the acid and base dissociation constants of the PE and NPs on the fraction of adsorbed monomers, $f_{ad}$. To this end, we systematically vary the values of $pK_{a}$ and $pK_{b}$ while keeping the salt concentrations fixed ($pIp=pIm=7$) and maintaining a constant pH of 7. To capture the combined influence of these parameters, we define the quantity $\Delta pK=\Delta pK_{a}+\Delta pK_{b}=(pK_{a}-pH)+(pK_{b}-pH)=pK_{a}+pK_{b}-2pH$. A positive $\Delta pK$ corresponds to a regime in which fewer surface sites on the NPs are protonated and fewer monomer units on the PE are deprotonated. Consequently, both the NPs and the PE bear reduced net charges, resulting in weaker electrostatic interactions and reduced adsorption. In contrast, a negative $\Delta pK$ indicates stronger acid/base dissociation, leading to more highly charged NPs and PE chains and, hence, stronger electrostatic attraction and greater adsorption. When $\Delta pK=0$, the system is effectively charge-balanced, representing an intermediate regime of interaction strength.
@@
Figure 5a shows that at $\Delta pK=0$, the net charges on both the NPs and the PE are balanced and relatively low, leading to weak electrostatic interactions. As a result, the corresponding fraction of adsorbed monomers, $f_{ad}$, remains low in both CR and CC simulations, with only a minor difference between the two cases (see Fig. 5b). As $\Delta pK$ becomes increasingly negative, the net charges on the NPs and PE grow, thereby strengthening the electrostatic attraction and enhancing adsorption. However, because the PE chain length is fixed at $N=40$, the total charge on the PE saturates beyond a certain threshold. Interestingly, the disparity between CR and CC simulations becomes most pronounced at intermediate $\Delta pK$ values. This can be attributed to the enhanced responsiveness of surface charge regulation under moderately strong electrostatic conditions. At very negative $\Delta pK$, although the electrostatic attraction is strong, the high adsorption in both cases narrows the difference between CR and CC again. Thus, we conclude that the distinction between CR and CC is minimal at $\Delta pK=0$, increases at intermediate values, and diminishes at large negative $\Delta pK$, where adsorption is near saturation in both regimes.
@@
Furthermore, we investigate the temporal evolution of the adsorbed monomer fraction, $f_{ad}$, at various $\Delta pK$ values for both CR and CC simulations. As shown in Fig. 6a, CR simulations exhibit a monotonic increase in the equilibrium value of $f_{ad}$ with increasingly negative $\Delta pK$, reflecting stronger electrostatic interactions and enhanced PE adsorption. A similar trend is observed in CC simulations (see Fig. 6b); however, the kinetics of adsorption differ markedly. Notably, the adsorption process proceeds significantly faster in the CR simulations, reaching equilibrium values approximately twice as quickly as in the CC case. This accelerated adsorption under charge-regulating conditions highlights the dynamic adaptability of surface charge, which facilitates rapid electrostatic accommodation and polymer attachment.
@@
Figure 6: Temporal evolution of adsorbed monomer fraction ($f_{ad}$) in (a) CR simulations and (b) CC simulations.
Finally, we investigate the electrostatic interaction between PE and NPs at a higher salt concentration ($pIp=pIm=3$), while keeping all other parameters identical to those used in Fig. 3. Simulation snapshots at $D=7\ell_{b}$ are depicted in Fig. 7a, revealing similar structural features to those observed under low salt conditions (see Fig. 3a). However, as illustrated in Fig. 7b, the charge on the NPs is nearly twice the magnitude and opposite in sign to that of the PE, due to the increased availability of free ions participating in surface ionization reactions. The impact of salt concentration on the interaggregate force is depicted in Fig. 7c.
In the CC simulations, polymer bridging persists, but at a shorter range of separation than in the low-salt scenario, which is consistent with previous findings Podgornik et al. (1995). In contrast, the CR simulations show no qualitative change in the force profile with increasing salt concentration, except for a noticeable downward shift in $F_{X}$. This reduction in net force arises from the increased population of counterions (primarily anions) required to neutralize the higher positive surface charge on the NPs. Overall, these observations suggest that, under CC conditions, NP-NP interactions are strongly PE-mediated at low salt concentrations but become progressively screened as the ionic strength increases. In contrast, under CR conditions, the interaggregate force remains dominated by counterions across salt concentrations, reflecting the dynamic charge regulation and strong ion-mediated screening intrinsic to these systems. Furthermore, we observe a significant increase in PE adsorption for both CR and CC simulations at high salt concentrations, as shown in Fig. 7d. Importantly, the difference in adsorption between CR and CC systems is significantly reduced compared to that seen under low-salt conditions (see Fig. 3d), which can be attributed to the higher surface charge on the NPs under these conditions. A similar convergence is evident in the radius of gyration (see Fig. 7e), where the disparity between CR and CC simulations diminishes with increasing salt concentration, indicating more compact PE conformations in both cases.
These results underscore that while the CR and CC systems show clear differences at intermediate separations–particularly under low ionic strength—such differences become less significant at larger separations and higher salt concentrations. In CR simulations, increased salt primarily leads to a suppression of interaggregate forces due to enhanced screening by counterions. In contrast, CC simulations remain sensitive to salt, with polymer-mediated bridging interactions being progressively diminished. Nevertheless, in both scenarios, higher salt promotes greater PE adsorption and a reduction in $R_{g}$, reflecting more compact and surface-associated polymer configurations.

### v2

@@ III Computation of total force between two nanoparticles
We now focus on the analytical form of the total force between two NPs due to bathing the PE solution. In the limit of CC, the interaction forces between charged colloids mediated by PEs have already been discussed in detail in Ref. Podgornik et al. (1995). However, we present a pedagogical derivation of these results to ensure that the paper is self-contained. To evaluate the total force acting between two NPs, we introduce a fictitious plane, $S$, of dimension $0.25\ell_{B}\times 80\ell_{B}\times 80\ell_{B}$, positioned at the midpoint between the two NPs (see Fig. 2)  Podgornik et al. (1995); Granfeldt et al. (1991). The system is divided into two sub-volumes, $V_{1}$ and $V_{2}$, corresponding to the left and right halves of the simulation domain, with $S$ acting as the dividing midplane. Given Ref. Podgornik et al. (1995), the total force ($F_{X}$) acting along the radius vector joining the two NPs can be cast as,
$\begin{split}F_{X}&=k_{B}T\int_{(S)}d^{2}\mathbf{r}\,\rho_{1}(\mathbf{r})\\ &\quad+\int_{(V_{1})}\int_{(V_{2})}d^{3}\mathbf{r_{1}}\,d^{3}\mathbf{r_{2}}\,\rho_{2}(\mathbf{r_{1}},\mathbf{r_{2}})\,\mathbf{f_{x}}(\mathbf{r_{1}},\mathbf{r_{2}})~.\end{split}$ (4)
@@
The first term of Eq. 4 represents the momentum exchange between $V_{1}$ and $V_{2}$ due to the movement of ions/particles across the midplane. It has the form of an ideal osmotic pressure Lovett and Baus (1992). The second term involves the microscopic force $\mathbf{f_{x}}(\mathbf{r_{1},r_{2}})$, which acts between a particle at position $\mathbf{r_{1}}\in V_{1}$ and another at $\mathbf{r_{2}}\in V_{2}$ projected along the direction normal to the dividing plane $S$. Here, $\rho_{1}(r)$ refers to the single particle density at the midplane and $\rho_{2}(\mathbf{r_{1}},\mathbf{r_{2}})$ is the two-particle correlation function for the system. This expression for force (Eq. 4) is related to the contact theorem in plan-parallel geometry Lovett and Baus (1992). The microscopic force $\mathbf{f_{x}}(\mathbf{r_{1},r_{2}})$ comprises several contributions: the electrostatic force between the polymer beads, the electrostatic force between the polymer beads and the two NPs, and the configurational (elastic) forces arising from bonded polymer segments whose connecting vector intersects the dividing plane $S$. As in Ref. Podgornik et al. (1995), we obtain the total force,
$F_{X}=F_{X}(\text{osm})+F_{X}(\text{cor})+F_{X}(\text{nn})+F_{X}(\text{bridge})~,$ (5)
@@
where $F_{X}(\text{osm})$ denotes the contribution from the osmotic force due to the momentum exchange mechanism across planes. The force component $F_{X}(\text{cor})$ is due to the correlation effect, and the contribution due to the electrostatic interaction between CR NPs is captured by $F_{X}(\text{nn})$. Since the polymer can act as a bridge between NPs when crossing mid-plane $S$, there is an additional term $F_{X}(\text{bridge})$ in the above Eq. 5.
Figure 2: Schematic of the system for the force calculation. A fictitious plane, $S$, of width $0.25\ell_{B}$ is placed at the midpoint, which divides the system into two halves, $V_{1}$ and $V_{2}$. The color scheme is the same as in Fig. 1.
@@
The osmotic force component, $F_{X}(\text{osm})$, is calculated using the relation $F_{X}(\text{osm})=kT<n>$, where <n><n> is the average number of particles inside the plane $S$. The correlation force component, $F_{X}(\text{cor})$, is obtained by evaluating the electrostatic interactions between (i) all mobile particles in subvolumes $V_{1}$ and $V_{2}$; (ii) all mobile particles in $V_{1}$ and the NP in $V_{2}$; and (iii) the NP in $V_{1}$ and all the mobile particle in $V_{2}$. $F_{X}(\text{nn})$ is direct electrostatic force between two NPs. We have utilized the Particle–Particle Particle–Mesh (PPPM) concept to compute electrostatic forces with a precision of $10^{-3}$. The bridging force, $F_{X}(\text{bridge})$, given by
$F_{X}(\text{bridge})=-K_{F}\left<|X_{i}-X_{i+1}|\right>\times n_{br}~,$ (6)
@@ IV Results and Discussion
In this section, we present numerical results based on the formulation described in the previous section. To elucidate the effects of charge regulation (CR) on the PE-NP interaction and NP-NP interaction, we compare the results of CR and constant charge (CC) simulations performed under the same conditions, as outlined in Sec. II.
@@
We begin with the case of a deionized solution, i.e., with $pIp=pIm=7$. The corresponding system parameters are $pH=7$, $pK_{a}=pK_{b}=5$, and the number of monomers is fixed at $N=40$. Figure 3a shows simulation snapshots of CR and CC systems at a fixed intersurface separation of $D=7\ell_{B}$. Notably, in the CR case, the PE is adsorbed onto the surface of one of the NPs (a localized state). In contrast, under CC conditions at the same separation, PE preferentially forms a bridge between two like-charged NPs, referred to here as the captured state. Within the CC approximation, bridging interactions occur due to the electrostatic attraction between oppositely charged PE chain and the fixed surface charges of NPs, creating a strong and persistent electrostatic attraction for the oppositely charged PE segments, leading to bridging. As a result, the PE becomes partially adsorbed onto both NPs, and the intervening segment forms a bridge that generates an entropic-elastic force between them. These observations are consistent with the previous studies Podgornik (2003); Podgornik and Ličer (2006). In contrast, under CR conditions, the surface charges of the NPs dynamically adjust in response to the local electrostatic environment. This flexibility allows one of the NPs to acquire a higher effective surface charge that strongly attracts the entire PE chain, resulting in its complete adsorption. This adaptive charge response suppresses polymer bridging and favors a localized state, thereby minimizing the system’s free energy through a combination of enhanced electrostatic attraction and configurational entropy gain. Further, we find that, under CR conditions, the PE is equally likely to adsorb onto either of the two NPs, reflecting the system’s symmetry.
@@
At equilibrium, the total charge on the PE and the adsorbing NPs is nearly equal in magnitude and opposite in sign for the CR parameters mentioned above, as illustrated in Fig. 3b. Figure 3c presents the total force ($F_{X}$) between the NPs as a function of their intersurface separation. The results show that the PE mediates an attractive interaction between like-charged NPs over a broad range of separations, consistent with previous findings for the CC case Podgornik et al. (1995). In contrast, for the CR case, attractive interactions are observed only at very short intersurface separations and diminish rapidly with increasing distance, becoming weakly repulsive around $D=7\ell_{B}$. This behavior results from the complete adsorption of the PE onto one of the NPs, effectively neutralizing its surface charge. As the separation increases, the dominant contribution to the total force arises from the osmotic pressure of mobile ions and monomers near the midplane, leading to a net repulsive interaction. Notably, charge regulation significantly alters the interaction landscape. As seen in Fig. 3c at $D=15\ell_{B}$, the magnitude of $F_{X}$ in the CR case differs by nearly a factor of six compared to that in CC. However, at very large separations, the total forces in both CR and CC converge. This is because, in the CC scenario, the PE also eventually undergoes complete adsorption onto an NP.
@@
Figure 3: Charge regulation (CR) effects on NPs–PE interactions at a low salt concentration ($pIp=pIm=7$) with PE bearing acidic groups ($pK_{a}=5$) and NPs coated with basic groups ($pK_{b}=5$). (a) Top: Snapshot of CR simulation at intersurface separation between colloids, $D=7\ell_{B}$. Red and white beads denote charged and neutral acid monomers of the PE, while blue and white sites represent charged and neutral base sites. Bottom: Simulation snapshot of a constant charge (CC) simulation at the same $D$. (b) The average charge on NPs and PE as a function of $D$. (c) The total force ($F_{X}$) between NPs for CR and CC as a function of $D$. (d) Adsorbed monomer fraction ($f_{ad}$), within $2\sigma$ of the colloid surface as a function of $D$. (e) The radius of gyration ($R_{g}$) of the PE as a function of $D$. The averages are calculated over 7 (CR) and 30 ($C$) equilibrated configurations.
@@
A quantitative characterization of the CR effect is provided by examining the fraction of adsorbed monomers, $f_{ad}$, defined as those located within $2\sigma$ of the NP surface (see Fig. 3d). At very small intersurface separations, monomers remain near the NP surfaces due to limited available space, resulting in similar values of $f_{ad}$ and $F_{X}$ for both CR and CC simulations. However, as $D$ increases beyond $7\ell_{B}$, the PE in CR simulations remains fully adsorbed onto one of the NP surfaces. As a consequence, attractive interactions are suppressed because they originate from a monomer that is stretched across the midplane. In contrast, for the CC case, $f_{ad}$ decreases with increasing $D$, indicating that the PE bridges the two NPs. As separation increases, monomers span a larger distance across the midplane, leading to stronger entropic-elastic attractions. Beyond a critical separation (around $15\ell_{B}$), the PE in CC simulations begins to adsorb onto a single NP surface, reducing the bridging and hence diminishing the attractive force. Even at very large separations, the CR model shows nearly twice the monomer adsorption compared to CC. This suggests that CR enhances the affinity of the PE for the surface, leading to stronger adhesion. A similar suppression of bridging at high surface charge has been observed experimentally. Popa et al. Popa et al. (2010) used colloidal-probe atomic force microscopy to show that once the surfaces reached or exceeded charge reversal through polyelectrolyte adsorption, bridging became rare despite high polymer coverage, consistent with our CR results. This difference is also reflected in the radius of gyration, $R_{g}$, shown in Fig. 3e, which remains consistently higher in CC simulations. The increased $R_{g}$ indicates a more extended polymer conformation due to weaker adsorption and stronger monomer-monomer repulsions. These findings indicate that CR significantly enhances PE adsorption onto NPs, in agreement with the recent studies on PE adsorption onto planar surfaces Yuan and Tanaka (2024).
@@
Next, we examine the influence of the length of the PE chain ($N$) on the adsorption behavior and the interaggregate force ($F_{X}$). Simulations were performed for $N=30$, 4040, and $50$ while keeping all other parameters identical to those used in Fig. 3. As shown in Fig. 4a, the length of the PE chain has a minimal impact on $F_{X}$ in CR simulations. This insensitivity can be attributed to the strong adsorption of the PE onto the surface of a single NP, as illustrated in Fig. 4b, which effectively neutralizes the NPs charge and suppresses long-range electrostatic interactions. Interestingly, Fig.4b shows that PE adsorption onto the NP surface monotonically increases with polymer chain length. This behavior is in stark contrast to previous observations for planar surfaces Yuan and Tanaka (2024), where adsorption tends to decrease with increasing chain length due to the entropic penalty associated with flattening long chains onto a surface. In our case, the curved geometry of the NP reduces the loss of conformational entropy during adsorption, allowing longer chains to maintain configurational flexibility while still forming multiple contacts with the surface. Additionally, the enhanced total charge of longer PEs increases their electrostatic attraction to the charged NP, further promoting adsorption. The CC model reveals a pronounced dependence of $F_{X}$ on the PE chain length. As depicted in Fig. 4c, longer chains lead to stronger attractive interactions at larger intersurface separations. This enhancement is driven by more effective bridging between the NPs, consistent with the observed decrease in PE adsorption onto individual NP surfaces (see Fig. 4d). Thus, while the interaggregate force in the CR case remains largely unaffected by the length of the PE because of complete adsorption, in the CC case the length dependence is strong because of persistent bridging interactions.
@@
Figure 4: Effect of PE chain length on interaggregate force and PE adsorption across different simulation models, under the same conditions as in Fig. 3. (a) The total force ($F_{X}$) between NPs in the CR simulation. (b) Adsorbed monomer fraction onto NPs in CR simulations. (c) $F_{X}$ between NPs in CC simulations. (d) Adsorbed monomer fraction onto NPs in CC simulation.
@@
Now, we investigate the effect of the acid and base dissociation constants of the PE and NPs on the fraction of adsorbed monomers, $f_{ad}$. To this end, we systematically vary the values of $pK_{a}$ and $pK_{b}$ while keeping the salt concentrations fixed ($pIp=pIm=7$) and maintaining a constant pH of 7. To capture the combined influence of these parameters, we define the quantity $\Delta pK=\Delta pK_{a}+\Delta pK_{b}=(pK_{a}-pH)+(pK_{b}-pH)=pK_{a}+pK_{b}-2pH$. A positive $\Delta pK$ corresponds to a regime in which a few surface sites on the NPs are protonated and fewer monomer units on the PE are deprotonated. Consequently, both the NPs and the PE bear reduced net charges, resulting in weaker electrostatic interactions and reduced adsorption. In contrast, a negative $\Delta pK$ indicates stronger acid/base dissociation, leading to more highly charged NPs and PE chains and, hence, stronger electrostatic attraction and greater adsorption. When $\Delta pK=0$, the system is effectively charge-balanced, representing an intermediate regime of interaction strength.
@@
Figure 5a exhibits that at $\Delta pK=0$, the net charges on both NPs and PE are balanced and relatively low, leading to weak electrostatic interactions. As a result, the corresponding fraction of adsorbed monomers, $f_{ad}$, remains low in both CR and CC simulations, with only a minor difference between the two cases as shown in Fig. 5b. As $\Delta pK$ becomes increasingly negative, the net charges on the NPs and PE grow, thereby strengthening the electrostatic attraction and enhancing adsorption. However, because the PE chain length is fixed at $N=40$, the total charge on the PE saturates beyond a certain threshold. Interestingly, the disparity between CR and CC simulations becomes most pronounced at intermediate $\Delta pK$ values. This can be attributed to the enhanced responsiveness of surface charge regulation under moderately strong electrostatic conditions. At a very negative $\Delta pK$, although the electrostatic attraction is strong, the high adsorption in both cases narrows the difference between CR and CC again. Thus, we conclude that the distinction between CR and CC is minimal at $\Delta pK=0$, increases at intermediate values, and diminishes at large negative $\Delta pK$, where adsorption is near saturation in both regimes.
@@
Furthermore, we investigate the temporal evolution of the adsorbed monomer fraction, $f_{ad}$, at various $\Delta pK$ values for both CR and CC simulations. As shown in Fig. 6a, CR simulations exhibit a monotonic increase in the equilibrium value of $f_{ad}$ with an increasing negative $\Delta pK$, reflecting stronger electrostatic interactions and improved PE adsorption. A similar trend is observed in CC simulations as evident in Fig. 6b; however, the kinetics of adsorption differ markedly. Notably, the adsorption process proceeds significantly faster in the CR simulations, reaching equilibrium values approximately twice as quickly as in the CC case. This accelerated adsorption under charge-regulating conditions highlights the dynamic adaptability of surface charge, which facilitates rapid electrostatic accommodation and polymer attachment.
@@
Figure 6: Temporal evolution of the adsorbed monomer fraction ($f_{ad}$) in (a) CR and (b) CC simulations. Blue, magenta, and metallic blue solid lines correspond to $\Delta pK=0$, $\Delta pK=-4$, and $\Delta pK=-8$, respectively.